OpenStudy (anonymous):

CHECK MY ANSWER WILL MEDAL. ( HOW MANY WAYS CAN TWO OUT OF 10 CHILDREN BE LINED UP, FOLLOWED BY LINING UP FOUR OUT OF 10 MEN, WHERE ORDER MATTERS, AND THEN INCLUDE 3 OUT OF 10 WOMEN WHERE ANY COMBINATION WILL WORK? my answer) 7,357,000

OpenStudy (mathmate):

10 children choose 2 = C(10,2) = 10!/(2!*8!)=36 10 men choose 4=C(10,4)=210 10 women choose 3 = C(10,3)=120 Multiply them together will give the number of ways to choose the 9 people (>1 million ways). Now multiply by the number of ways you can line up the 9 people to get the final answer.

OpenStudy (anonymous):

Yah ^^ plzz fan me and medal @sar12389 i was going to say same thing thx :)

OpenStudy (mathmate):

:)

OpenStudy (anonymous):

Thx @mathmate